If MPC = 0.75 (and there are no income taxes) when G increases by 100, then the IS curve for any given interest rate shifts to the right by:

A) 100.

B) 200.

C) 300.

D) 400.
Answer #1

Ans: D ) 400

Explanation:

Multiplier = 1 / ( 1 - MPC )

= 1 / ( 1 - 0.75 )

= 1 / 0.25

= 4

If MPC = 0.75 (and there are no income taxes) when G increases by 100, then the IS curve for any given interest rate shifts to the right by ;

= Government spending * multiplier

=100 * 4

= 400